If accelerating potential increases from20kv to 80kv in an electron microscope, its resolving power R would change to

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Sayam
 
\lambda = h / P = (h / \small \sqrt2me)(1 / \small \sqrtv)
R.P. \small \alpha 1 / \lambda  \small \alpha  \small \sqrtv
R / R’ = \small \sqrt(20/ 80) = 1/ 2
R’ = 2R
hope it helps
Regards
Arun (askIITians forum expert)